OpenStudy (anonymous):

Use your knowledge of the derivative to compute the value of the limit given below: \lim_{h \rightarrow 0} \frac{\sqrt[1]{(x + h)^{3}} - \sqrt[1]{x^{3}}}{h} . The derivative is \frac{dy}{dx}= .

OpenStudy (zarkon):

\sqrt[1]{(x + h)^{3}} ... \[\sqrt[1]{(x + h)^{3}} \] what do you mean by this?

OpenStudy (anonymous):

see attached

OpenStudy (zarkon):

that is either a typo or just weird use of notation...

OpenStudy (zarkon):

\[\sqrt[n]{x}=x^{1/n}\] so \[\sqrt[1]{x}=x^{1/1}=x\]

OpenStudy (zarkon):

so you could try \[3x^2\]

OpenStudy (zarkon):

I'm betting the problem is a typo

OpenStudy (anonymous):

thank 3x^2 is correct, ur a genius - i have been trying to figure this out all night

OpenStudy (zarkon):

very poor use of notation